OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y =4/x^2 + 9

OpenStudy (anonymous):

nevermind i did it

OpenStudy (anonymous):

f(x) is 4/x^2 and g(x) is 9

OpenStudy (anonymous):

f(x) = 4/x +9 g(x) = X^2

OpenStudy (anonymous):

hmm it's multiple choice and i do not see that there